# Prediction of Cellular Malignancy Using Electrical Impedance Signatures and Supervised Machine Learning

### v1

@@ 1 Introduction
According to the Centers for Disease Control and Prevention (CDC) and the National Cancer Institute, approximately 614,000 cancer-related deaths were reported in the United States in 2023. Additionally, an estimated 1.9 million and 2.0 million new cancer cases were reported in 2022 and 2025, respectively. Several factors influence cancer prognosis including cancer type, (ii) stage at diagnosis, (iii) cellular characteristics, and (iv) histological grade, among others [Rakha2010]. Notably, changes in the electrical properties and biochemical composition of cells have emerged as significant indicators for disease diagnosis and progression [Hossain2020]. Early detection is strongly correlated with improved survival outcomes, underscoring the importance of timely diagnosis and monitoring to reduce mortality and expand treatment options [McPhail2015],[Muller2018],[vandenBergh2015].
@@
One of the key advantages of electrical characterization of cells is its non-invasive and label-free nature [NgocLe2019]. Unlike fluorescence -based techniques, dielectric measurement does not require staining or labeling, which can potentially alter cellular behavior. Moreover, cells remain viable after measurement, enabling further downstream analysis or long-term studies if needed. Another notable advantage, driven by advancements in medical technology, is that real-time dynamic electrical measurements enable continuous cell monitoring and facilitate high-throughput screening.
@@
Cheung et al. (2005) developed a microfabricated impedance spectroscopy flow cytometer for the rapid dielectric characterization of cells across a range of frequencies [Cheung2005]. They utilized electrical features such as amplitude, phase, and opacity to actively discriminate between cell types without the use of molecular markers. Although the study was limited to red blood cells (RBC), the approach holds potential for broader applications, including the measurements of cell conductivity and capacitance for early-stage apoptosis or cancer diagnostics.
@@
Similarly, Gawad et al. (2001) introduced a cytological tool capable of high-speed cell counting and separation, operating at a sampling rate of 100 samples per second [Gawad2001]. Their chip-based flow cytometer employed a differential pair of microelectrodes to measure impedance over a frequency range of 100 kHz to 15 MHz. This system can be extended to distinguish between normal and malignant cells based on differences in bioimpedance or conductivity.
@@
Our previous works focused on electrical impedance and transmembrane potential as biophysical indicators for cancer cell diagnosis [Hossain2021],[Hossain2020]. The study revealed that cancer cells exhibit higher electrical conductivity and greater dielectric loss compared to their healthier counterparts. Additionally, cancer cells tend to be depolarized, often displaying elevated intracellular sodium ion concentrations despite similar external ionic conditions [Kunzelmann2005],[Bortner2014],[Binggeli1986]. For instance, ovarian tumor cells have reported to possess a transmembrane potential of approximately -5 mV, while MCF -7 breast cancer cells range between -37 to -38 mV , and MDA- MB-468 cells exhibit a potential around -30 mV [Mahmoud2023]. In our previous works, we also investigated that both the bioimpedance and optical properties of breast cancer cells can be used to enhance the precision of malignant cell identification [Hossain2020],[Hossain2021],[HossainHossain2021]. Electrical measurements were conducted in the microwave frequency range, while the optical characteristics of normal and cancer cells were evaluated in the visible and near- infrared spectral (NIR) regions. Similarly, Alfano et al. conducted several pioneering studies utilizing optical techniques to distinguish precancerous and cancerous tissues based on their characteristic absorption wavelength [Alfano2016],[Alfano2013],[Alfano2005]. However, despite their potential, optical diagnostic techniques face significant challenges, including :(i) limited spatial resolution, (ii) prolonged acquisition time, and (iii) reduced sensitivity in complex biological environments [Licha2005].
@@
Measurable bioelectrical differences between cancerous and non-cancerous cells provide a valuable feature space for computational classification. By leveraging machine learning algorithms, such as Support Vector Machine (SVM), Random Forest (RF), Deep Neural Network (DNN) or K-Nearest Neighbor (KNN), it is possible to learn patterns in electrical impedance, dielectric loss and transmembrane potential that distinguish malignant cells from healthy ones [Biau2016],[Rigatti2017],[Guo2003],[Hossain2024Authorea],[Suthaharan2016]. These models can be trained on labeled datasets derived from impedance cytometry or microelectrode array recordings, enabling real time non-invasive cancer diagnostics with high accuracy. Moreover, the ability of machine learning to capture complex, non-linear relationships enhance the diagnostic potential beyond traditional threshold-based methods. Fig.1 shows the schematic of the sequence of steps for the proposed model discussed earlier. Bioelectrical measurements -such as impedance, conductivity, and dielectric loss -are acquired from test samples using microelectrodes. These features are then input into a pretrained machine learning classifier for binary classification, distinguishing between malignant and healthy cells based on their electrical signatures.
@@ 2.1 Bioelectrical Properties and Cole-Cole plot
Cancer progression is accompanied by distinct biophysical and biochemical changes at the cellular level, many of which affect the structural and electrical properties of the cell membrane and cytoplasm [Hossain2020],[Hossain2021]. As a result electrical characterization techniques have emerged as a valuable tool in cancer research, offering non-invasive label-free methods to distinguish between normal, premalignant and malignant cells. Among these techniques, dielectrophoresis (DEP) has shown promise in detecting subtle differences in detecting dielectric properties such as permittivity, conductivity and membrane capacitance [Pethig2010],[Pohl1971],[Nakano2013],[Hossain2023],[Hossain2020],[Lapizco2007]. These properties which are frequency dependent can be modeled using frameworks like the Cole-Cole equation to reveal patterns that correlate with malignancy. The ability to extract quantitative dielectric signatures makes the approaches particularly suitable for early detection, classification, and monitoring of cancer progression in vitro and in vivo settings.
@@
A biological cell can be electrically modeled as a combination of resistive and capacitive elements [Brosseau2021],[Gowrishankar2003],[Ellappan2005]. The cytoplasm, rich in ions, is represented by an internal resistance, Ri reflecting its conductive properties. In contrast, the cell membrane, primarily composed of a lipid bilayer, acts as a dielectric barrier and is modeled as a parallel combination of resistance and capacitance. This RC configuration accounts for the membrane’s ability to store and partially block charge flow. Importantly, the overall importance of the cell is frequency-dependent: at low frequencies, the capacitive nature of the membrane dominates, impending current flow; while at higher frequencies, the capacitive reactance decreases, allowing current to pass through resistive pathways more easily [Trainito2015],[Mahesh2020],[Karmakar2024].
@@
Giannoukis and Min (2014) explored the mathematical and physical modeling of the dynamic electrical bioimpedance of cells [Giannoukos2014]. Fig. 2(a) illustrates the Frickle-Morse equivalent circuit model representing a biological cell. Fig. 2(b) and (c) demonstrate the variation in electric field distribution within normal and cancerous cells, representing how differences in dielectric properties influence the field patterns. The increased permittivity of cancerous cells alters the dielectric response, resulting in more concentrated electric field distribution compared to non-tumorous.
Figure 2: (a) The Frickle-Morse equivalent circuit model used for impedance extraction. (b-c) Comparison of electric field lines between healthy and malignant cell morphologies.
@@
The Cole- Cole equation is the relaxation model that is used to describe the dielectric relaxation, and the model can also be adapted to compare the separation in dielectric properties of cancerous and non-cancerous breast cancer cells [Cole1941]. Fig. 3 shows the Cole-Cole plot illustrating the separation in dielectric properties between cancerous and non-tumorous breast cell lines. The plot generated from dielectric spectroscopy data of various breast cell lines (including both cancerous and non-tumorigenic types), demonstrates a clear clustering pattern. The spatial distribution of the complex permittivity values suggests the existence of a separating hyperplane, enabling potential binary classification based on the dielectric properties. Further quantitative analysis (e.g. linear discrimination analysis or SVM classification) is warranted to validate the robustness and generalizability of this observed separation and is the focus of the following section.
@@
Similar studies by Roberts et al. (2005) investigated the progression of mouse ovarian surface epithelial (MOSE) cells [Roberts2005]. This work analyzed cellular and molecular changes across the different stages, from pre-malignant, non-tumorigenic states to highly aggressive malignant phenotypes. The model includes four transitional states including MOSE-E (early preneoplastic stage) and MOSE-L (the most aggressive stage) with high tumorigenic potential.
In addition, to dielectric spectroscopy, studies on dielectrophoresis (DEP) – a phenomenon in which a force is exerted on dielectric particle in a non-uniform electric field has also been employed to investigate the different stages of MOSE cells [Trainito2019]. The results indicate that as malignancy progresses the membrane capacitance decreases from approximately 26 mF/m2 to 15 mF/m2. This decline highlights the difference in dielectric properties and membrane composition between malignant and non-tumorous cells, reinforcing the potential of electrical characterization methods in distinguishing cancer progression.
While dielectric properties provide a promising biophysical basis for distinguishing between malignant and non-malignant cells, the complexity and variability of the measured parameters necessitate robust computational tools for accurate classification. Artificial Intelligence (AI) and Machine Learning (ML) algorithms have been increasingly adopted to analyze bioelectrical data. These models can learn patterns from features such as membrane capacitance, conductivity, and impedance spectra to reliably differentiate between healthy and cancerous cells. Among various ML techniques, models such as Random Forest (RF), Support Vector Machine (SVM) and K-Nearest Neighbor (KNN), and Neural Network have shown promising results in biomedical classification tasks due to their ability to handle non-linear data and high dimension feature space [Breiman2001],[Liu2012],[Peterson2009],[Kramer2013],[Hearst1998],[Noble2006],[Hossain2021].

### v2

@@ 1 Introduction
According to the Centers for Disease Control and Prevention (CDC) and the National Cancer Institute, approximately 614,000 cancer-related deaths were reported in the United States in 2023. Additionally, an estimated 1.9 million and 2.0 million new cancer cases were reported in 2022 and 2025, respectively. Several factors influence cancer prognosis including cancer type, (ii) stage at diagnosis, (iii) cellular characteristics, and (iv) histological grade, among others [1]. Notably, changes in the electrical properties and biochemical composition of cells have emerged as significant indicators for disease diagnosis and progression [2]. Early detection is strongly correlated with improved survival outcomes, underscoring the importance of timely diagnosis and monitoring to reduce mortality and expand treatment options [3],[4],[5].
@@
One of the key advantages of electrical characterization of cells is its non-invasive and label-free nature [6]. Unlike fluorescence -based techniques, dielectric measurement does not require staining or labeling, which can potentially alter cellular behavior. Moreover, cells remain viable after measurement, enabling further downstream analysis or long-term studies if needed. Another notable advantage, driven by advancements in medical technology, is that real-time dynamic electrical measurements enable continuous cell monitoring and facilitate high-throughput screening.
@@
Cheung et al. (2005) developed a microfabricated impedance spectroscopy flow cytometer for the rapid dielectric characterization of cells across a range of frequencies [7]. They utilized electrical features such as amplitude, phase, and opacity to actively discriminate between cell types without the use of molecular markers. Although the study was limited to red blood cells (RBC), the approach holds potential for broader applications, including the measurements of cell conductivity and capacitance for early-stage apoptosis or cancer diagnostics.
@@
Similarly, Gawad et al. (2001) introduced a cytological tool capable of high-speed cell counting and separation, operating at a sampling rate of 100 samples per second [8]. Their chip-based flow cytometer employed a differential pair of microelectrodes to measure impedance over a frequency range of 100 kHz to 15 MHz. This system can be extended to distinguish between normal and malignant cells based on differences in bioimpedance or conductivity.
@@
Our previous works focused on electrical impedance and transmembrane potential as biophysical indicators for cancer cell diagnosis [9],[2]. The study revealed that cancer cells exhibit higher electrical conductivity and greater dielectric loss compared to their healthier counterparts. Additionally, cancer cells tend to be depolarized, often displaying elevated intracellular sodium ion concentrations despite similar external ionic conditions [10],[11],[12]. For instance, ovarian tumor cells have reported to possess a transmembrane potential of approximately -5 mV, while MCF -7 breast cancer cells range between -37 to -38 mV , and MDA- MB-468 cells exhibit a potential around -30 mV [13]. In our previous works, we also investigated that both the bioimpedance and optical properties of breast cancer cells can be used to enhance the precision of malignant cell identification [2],[9],[14]. Electrical measurements were conducted in the microwave frequency range, while the optical characteristics of normal and cancer cells were evaluated in the visible and near- infrared spectral (NIR) regions. Similarly, Alfano et al. conducted several pioneering studies utilizing optical techniques to distinguish precancerous and cancerous tissues based on their characteristic absorption wavelength [15],[16],[17]. However, despite their potential, optical diagnostic techniques face significant challenges, including :(i) limited spatial resolution, (ii) prolonged acquisition time, and (iii) reduced sensitivity in complex biological environments [18].
@@
Measurable bioelectrical differences between cancerous and non-cancerous cells provide a valuable feature space for computational classification. By leveraging machine learning algorithms, such as Support Vector Machine (SVM), Random Forest (RF), Deep Neural Network (DNN) or K-Nearest Neighbor (KNN), it is possible to learn patterns in electrical impedance, dielectric loss and transmembrane potential that distinguish malignant cells from healthy ones [19],[20],[21],[22],[23]. These models can be trained on labeled datasets derived from impedance cytometry or microelectrode array recordings, enabling real time non-invasive cancer diagnostics with high accuracy. Moreover, the ability of machine learning to capture complex, non-linear relationships enhance the diagnostic potential beyond traditional threshold-based methods. Fig.1 shows the schematic of the sequence of steps for the proposed model discussed earlier. Bioelectrical measurements -such as impedance, conductivity, and dielectric loss -are acquired from test samples using microelectrodes. These features are then input into a pretrained machine learning classifier for binary classification, distinguishing between malignant and healthy cells based on their electrical signatures.
@@ 2.1 Bioelectrical Properties and Cole-Cole plot
Cancer progression is accompanied by distinct biophysical and biochemical changes at the cellular level, many of which affect the structural and electrical properties of the cell membrane and cytoplasm [2],[9]. As a result electrical characterization techniques have emerged as a valuable tool in cancer research, offering non-invasive label-free methods to distinguish between normal, premalignant and malignant cells. Among these techniques, dielectrophoresis (DEP) has shown promise in detecting subtle differences in detecting dielectric properties such as permittivity, conductivity and membrane capacitance [24],[25],[26],[27],[2],[28]. These properties which are frequency dependent can be modeled using frameworks like the Cole-Cole equation to reveal patterns that correlate with malignancy. The ability to extract quantitative dielectric signatures makes the approaches particularly suitable for early detection, classification, and monitoring of cancer progression in vitro and in vivo settings.
Fig.2 (a-f) compares three homogeneous baseline spherical cell models representing healthy, benign, and cancerous cell models to examine the influence of bulk dielectric properties on the internal electric field response. Fig. 2(a-c) shows the quantitative distribution of the scalar electric potential on the surface of the dielectric sphere for each cell type. The surface potential depends on: (i) the relative permittivity of the bulk cell, (ii) the magnitude of the externally applied electric field, and (iii) angular orientation with respect to the applied field direction.
Mathematically, the internal electric field ($E_{in}$) and surface potential ($\Phi$) of a homogeneous dielectric sphere subjected to a uniform external electric field are given by equations (1) and (2): [29],[30]:
$E_{in}=\left(\frac{3}{\epsilon_{r}+2}\right)E_{o}$ (1)
$\Phi(r,\theta)=-E_{in}R\cos\theta$ (2)
where $E_{in}$ is the uniform internal electric field inside the dielectric sphere, $E_{o}$ is the applied external field, $R$ is the radius of the cell, $\epsilon_{r}$ is the relative permittivity of the bulk medium, and $\theta$ is the polar angle measured from the field direction.
Relative permittivity values of 7.2, 18.3 and 59.15 were used to represent healthy, benign and cancerous cells, respectively. As the relative permittivity increases, a reduction in the surface electric potential magnitude is observed, indicating enhanced dielectric screening with the cell.
Figure 2: Quantitative simulation models comparing three homogeneous baseline cell representations: (a) healthy, (b) benign, and (c) cancerous, illustrating the influence of bulk dielectric properties on the internal electric field. The corresponding electric field magnitude distributions inside the dielectric sphere are shown for (d) healthy, (e) benign, and (iii) cancerous homogeneous cell.
Fig. 2(d–f) illustrates the corresponding electric field magnitude distribution inside the homogeneous spheres calculated using equation (3):
$|E|=\sqrt{E_{x}^{2}+E_{y}^{2}+E_{z}^{2}}$ (3)
where $E_{x}$, $E_{y}$, and $E_{z}$ are the Cartesian components of the electric field.
The electric potential exhibits a dipolar distribution characteristic of a polarized dielectric sphere, while the electric field magnitude remains spatially uniform within each homogeneous model. This behavior is consistent with the analytical solution of Laplace’s equation for a dielectric sphere embedded in a uniform electric field. The homogeneous model isolates bulk permittivity-driven field screening effects, providing a useful baseline for comparison with impedance-based measurements relevant to cellular diagnosis. However, real biological cells possess a thin insulating membrane and exhibit frequency-dependent conductivity and interfacial polarization effects, which are not captured in this simplified representation.
@@
A biological cell can be electrically modeled as a combination of resistive and capacitive elements [31],[32],[33]. The cytoplasm, rich in ions, is represented by an internal resistance, Ri reflecting its conductive properties. In contrast, the cell membrane, primarily composed of a lipid bilayer, acts as a dielectric barrier and is modeled as a parallel combination of resistance and capacitance. This RC configuration accounts for the membrane’s ability to store and partially block charge flow. Importantly, the overall importance of the cell is frequency-dependent: at low frequencies, the capacitive nature of the membrane dominates, impending current flow; while at higher frequencies, the capacitive reactance decreases, allowing current to pass through resistive pathways more easily [34],[35],[36].
Iqbal et al. (2019) presented an equivalent electrical model of an isolated single biological cell composed of resistive and capacitive elements [37]. In this model, the cell is represented by distinct compartments corresponding to the cytoplasm, nucleus and cell membrane. The capacitive components arise primarily from the charge-separation layers such as cell membrane and nuclear envelope, while the cytoplasm and nucleus are modeled as lossy conductive media and therefore contribute predominantly resistive components.
Electrical impedance measurements of biological cells reflect the combined contribution of resistive and capacitive elements. Effective capacitance is governed by the relative permittivity of cellular constituents, which is inherently frequency-dependent and influences the measured impedance response. Cancerous cells are reported to exhibit higher relative permittivity compared to healthy cells, leading to enhanced capacitive effects. These dielectric contrasts form the physical basis for impedance-based differentiation of cellular states.
@@
Giannoukis and Min (2014) explored the mathematical and physical modeling of the dynamic electrical bioimpedance of cells [38]. Fig. 3(a) illustrates the Frickle-Morse equivalent circuit model representing a biological cell. Fig.3 (b) and (c) demonstrate the variation in electric field distribution within normal and cancerous cells, representing how differences in dielectric properties influence the field patterns. The increased permittivity of cancerous cells alters the dielectric response, resulting in more concentrated electric field distribution compared to non-tumorous.
Figure 3: (a): Illustration of the Fricke-Morse equivalent circuit model representing a biological cell. (b) and (c): Variation in electric field distribution highlighting difference between normal cells and cells with higher effective permittivity (e.g. cancerous cells). (d) Schematic of impedance measurements using planar gold electrodes immersed in a conductive electrolyte medium, (e) Extended equivalent circuit model incorporating resistive and capacitive elements representing the membrane, cytoplasm and nucleus. (f) Frequency-dependent electrical behavior of a cell, illustrating current confinement at lower frequencies and membrane penetration at higher frequencies.
Fig. 3(d) illustrates the schematic of an impedance-based cell analysis technique, in which a small AC excitation current is applied across microelectrodes immersed in a conductive saline medium containing suspended cells. The presence of suspended cells perturbs the resulting current flow due to contrasts in their frequency-dependent dielectric properties, enabling the extraction of electrical impedance in a label free and real-time manner. This approach has been widely employed for dynamic cell analysis and characterization [37],[39],[40],[41]. Under the assumption of a dilute cell suspension and a simplified single-shell electrical representation, the effective impedance of the cell-electrolyte system can be approximated using equation (4)[39].
$Z_{mix}=\frac{R_{e}(1+j\omega R_{i}C_{m})}{j\omega R_{e}C_{m}+(1+j\omega R_{i}C_{m})(1+j\omega R_{e}C_{m})}$ (4)
where $R_{e}$ is the resistance of the electrolyte solution, $\omega$ is the angular frequency, $R_{i}$ is the cytoplasmic resistance, and $C_{m}$ is the membrane capacitance.
The frequency-depending dielectric contrast between the cell membrane and intracellular component leads to a measurable variation in current flow, from which the impedance spectrum can be extracted [41]. In Electric cell-substrate impedance sensing (ECIS), impedance is defined as the ratio of the measured voltage to the applied current across the electrodes over a range of AC frequencies[42]. Fig. 3 (e) presents a complex equivalent circuit model of a biological cell inspired by the Fricke-Morse single-shell model. The cytoplasm, owing to its intracellular conductivity is represented by an intracellular resistance $R_{i}$ while both the cell membrane and nucleus are modeled using parallel resistive and capacitive elements, characterized by membrane capacitance $C_{m}$ and nuclear capacitance $C_{n}$ , respectively. These capacitive components govern the frequency-dependent electrical behavior of the cell through capacitive coupling, as illustrated in Fig. 3(f). The capacitive reactance of the cell membrane is given by equation (5):
$X_{c}=\frac{1}{2\pi fC_{m}}$ (5)
At higher frequencies, the membrane reactance $X_{\text{c}}$ decreases, allowing electric current to penetrate the cell interior, whereas at lower frequencies the increased reactance inhibits the current flow through the membrane, confining it primarily to the extracellular region.
Several impedance-based techniques have been developed to protect the electrical properties of biological cells, each differing in electrode configuration, frequency range, sensitivity, and application domain. These methods exploit the frequency-dependent dielectric response of cell membranes and intracellular components to extract distinctive electrical impedance signatures associated with cellular morphology and physiological state. Table LABEL:tab:model_comparison1summarizes representative impedance measurements techniques reported in the literature, highlighting the cell types studied, key electrical parameters extracted, and their reported applications in tumor cell characterization.
Table 1: Comparison of impedance-based measurement techniques and reported electrical parameters of tumor cells from the literature
Technique Method Application Study Freq. Results Ref.
[…]
Electrical Impedance Spectroscopy (EIS) Analyzes complex impedance spectrum using equivalent circuit models over a broad range. Label-free characterization of dielectric properties and cell discrimination. SW403, Jurkat, and THP-1 10 Hz–1 MHz $R_{ct}$ (SW403/Jurkat): 2250 $\Omega$; THP-1: 2000 $\Omega$ [46]
Microfluidic cell impedance analysis Measures voltage/impedance changes as single cells flow through microchannels under AC excitation. Label-free discrimination based on size, membrane, and conductivity. A549 tumor cells, RBC, and WBC 0.1–10 MHz A549: 4.2–79.9 mV; RBC: 2.4–19 mV; WBC: 1.7–10.5 mV [47]
@@
The Cole- Cole equation is the relaxation model that is used to describe the dielectric relaxation, and the model can also be adapted to compare the separation in dielectric properties of cancerous and non-cancerous breast cancer cells [48]. Fig. 3 shows the Cole-Cole plot illustrating the separation in dielectric properties between cancerous and non-tumorous breast cell lines. The plot generated from dielectric spectroscopy data of various breast cell lines (including both cancerous and non-tumorigenic types), demonstrates a clear clustering pattern. The spatial distribution of the complex permittivity values suggests the existence of a separating hyperplane, enabling potential binary classification based on the dielectric properties. Further quantitative analysis (e.g. linear discrimination analysis or SVM classification) is warranted to validate the robustness and generalizability of this observed separation and is the focus of the following section.
@@
Similar studies by Roberts et al. (2005) investigated the progression of mouse ovarian surface epithelial (MOSE) cells [49]. This work analyzed cellular and molecular changes across the different stages, from pre-malignant, non-tumorigenic states to highly aggressive malignant phenotypes. The model includes four transitional states including MOSE-E (early preneoplastic stage) and MOSE-L (the most aggressive stage) with high tumorigenic potential.
In addition, to dielectric spectroscopy, studies on dielectrophoresis (DEP) – a phenomenon in which a force is exerted on dielectric particle in a non-uniform electric field has also been employed to investigate the different stages of MOSE cells [50]. The results indicate that as malignancy progresses the membrane capacitance decreases from approximately 26 mF/m2 to 15 mF/m2. This decline highlights the difference in dielectric properties and membrane composition between malignant and non-tumorous cells, reinforcing the potential of electrical characterization methods in distinguishing cancer progression.
While dielectric properties provide a promising biophysical basis for distinguishing between malignant and non-malignant cells, the complexity and variability of the measured parameters necessitate robust computational tools for accurate classification. Artificial Intelligence (AI) and Machine Learning (ML) algorithms have been increasingly adopted to analyze bioelectrical data. These models can learn patterns from features such as membrane capacitance, conductivity, and impedance spectra to reliably differentiate between healthy and cancerous cells. Among various ML techniques, models such as Random Forest (RF), Support Vector Machine (SVM) and K-Nearest Neighbor (KNN), and Neural Network have shown promising results in biomedical classification tasks due to their ability to handle non-linear data and high dimension feature space [51],[52],[53],[54],[55],[56],[9].